A car travels at a consistent speed of 30 mph. If the driver only travels a distance of 15,840 feet, how long will it take the driver, in minutes?
Answer by rothauserc(4718)

1 mile = 5,280 feet
:
15840/5280 = 3 miles
:
rate * time = distance
:
30 * time = 3
:
time = 3/30 = 1/10 of an hour
:
*********************************************
It will take the driver 60 * 1/10 = 6 minutes
*********************************************
: